# Global-Local Interface with Selective Direct and Singularity-Avoiding Motion Mapping for Intuitive Teleoperation

### v2

@@ III Direct Mapping of Local Component Motion
The global component of the Direct Mapping implementation covers the full six DoFs in the Pieper structure slave robot [33, 34, 35]. Activating the global component places the slave system in joint control mode, replicating the scaled replica’s joint positions within the limits of the motor capacities. Conversely, activating the local component switches the slave system to end-effector Cartesian control, where it tracks the motion of the haptic device.
To facilitate precise control of the end-effector, the slave system is designed to track only a portion of the local component motion. Denote the slave end-effector pose with $(\bm{p},\bm{R})$ where $\bm{p}\in\mathbb{R}^{3}$ is the position and $\bm{R}\in\mathbb{R}^{3\times 3}$ is the orientation represented by a rotational matrix. Let $\bm{q}$ be the joint poses of the slave robot, then we have $\bm{q}=\bm{F}(\bm{p},\bm{R})$ where $\bm{F}$ is the inverse kinematics of Pieper kinematics structure, which can be efficiently computed in closed form [32].
Every time the local component is activated, its starting pose $(\bm{p}^{m}_{0},\bm{R}^{m}_{0})$ as well as the slave end-effector’s starting pose $(\bm{p}^{s}_{0},\bm{R}^{s}_{0})$ will be recorded. Suppose at a certain moment $t$, the local component moves to pose $(\bm{p}^{m}_{t},\bm{R}^{m}_{t})$, we first compute the linear and rotational displacements since activation.
$\begin{split}\bm{dp}_{t}^{m}&=\bm{p}_{t}^{m}-\bm{p}_{0}^{m}\\ (\bm{l}^{m}_{t},\theta^{m}_{t})&=\bm{G}(\bm{dR}^{m}_{t})=\bm{G}(\bm{R}^{m,T}_{0}R^{m}_{t})\end{split}$ (1)
where the superscript $T$ stands for transpose and $\bm{G}$ maps the rotational matrix to axis-angle format. $\bm{l}_{t}$ is the normalized rotational axis direction and $\theta_{t}$ is the angle.
Given prespecified linear and rotational scaling factors $\alpha_{l},\alpha_{r}\in(0,1]$, the slave system will move its end-effector pose towards the scaled linear and rotational displacements of the haptic device. Mathematically, the target slave pose at moment $t$ is
$\begin{split}\bm{p}_{t}^{s}&=\bm{p}_{0}^{s}+\alpha_{l}\bm{dp}_{t}^{m}\\ \bm{R}_{t}^{s}&=\bm{R}_{0}^{s}\bm{G}^{-1}(\bm{l}^{m}_{t},\alpha_{r}\theta^{m}_{t})\end{split}$ (2)
with $\bm{G}^{-1}$ being the mapping that gets the rotational matrix from the axis-angle interpretation.
With the generated end-effector pose, the target joint poses of the slave robot at moment $t$ will be
$\bm{q}^{s}_{t}=\bm{F}(\bm{p}^{s}_{t},\bm{R}^{s}_{t})$ (3)
IV Singularity-Avoiding Mapping of Local Component Motion
Despite the preservation of full degrees of freedom of the end-effector, the aforementioned direct motion mapping has two major limitations.
  • It is particularly vulnerable against manipulator singularities, especially the wrist singularity in Pieper kinematics, which can happen everywhere inside the manipulator reachable workspace [36]. This demands human operators to pay extra attention on the slave joint poses other than the end-effector pose.
  • When the starting orientation of the local component and the slave end-effector have an obvious misalignment, controlling orientation using direct mapping can be overwhelming for human operators. This is because it is difficult for human operators to predict the slave end-effector’s orientation change given the motion of the local component. Additionally, they cannot intuitively discern which joints will be actuated and how much they will move.
This section focuses on the scenario when the slave system has a Pieper kinematic structure and proposes a practical local component motion mapping method to alleviate these limitations.
IV-A Pieper Kinematics and Translation-Rotation Decoupling
Pieper kinematics structure, also known as the 321 kinematics structure, is widely adopted in 6-joint industrial robot arms for its closed-form inverse kinematics. Fig. 3 is an illustration of the three most important characteristics of Pieper kinematics structure using AgileX Piper as an example.
  • The base joint (J1) and the first shoulder joint (J2) intersect and are perpendicular.
  • The two shoulder joints (J2 and J3) are parallel.
  • The three wrist joints (J4, J5 and J6) intersects at the same point $Q$, usually arranged in an intrinsic X-Y-X Euler Angles order.
Figure 3: (A) Pieper Kinematics Structure using Agile X Piper as an example. (B) $q_{5}=0^{\circ}$ wrist singularity, which is a typical manipulation pose and is also the singularity type we are focusing on avoiding.
It can be observed that the intersection point $Q$ is relatively close to the end-effector tip and its position is only affected by the poses of the base and shoulder joints. In addition, since the three wrist joints formulate an Euler angle combinations, they can dexterously adjust the end-effector orientation in a nonzero-measure set without actuating J1, J2 and J3, regardless of the Cartesian position of $Q$ [37]. With this observation, the following amendments are made to the Direct Mapping realization.
  • Translation of the local component now controls the linear motion of the intersection point $Q$, $\bm{p}_{t}^{Q}$, not the end-effector’s linear motion $\bm{p}_{t}^{s}$. Therefore, only J1, J2 and J3 will be actuated.
  • Rotation of the local component still controls the slave end-effector’s orientation $\bm{R}_{t}^{s}$, but only J4, J5 and J6 will be actuated.
Denote the starting pose of point $Q$ as $(\bm{p}_{0}^{Q},\bm{R}_{0}^{Q})$, the target position of point $Q$ at moment $t$ will be
$\bm{p}_{t}^{Q}=\bm{p}_{0}^{Q}+\alpha_{l}\bm{dp}_{t}^{m}$ (4)
The joint poses are now solved as
$\begin{split}\begin{bmatrix}q_{1,t}^{s},q_{2,t}^{s},q_{3,t}^{s}\end{bmatrix}^{T}&=\bm{F}_{123}(\bm{p}_{t}^{Q})\\ \begin{bmatrix}q_{4,t}^{s},q_{5,t}^{s},q_{6,t}^{s}\end{bmatrix}^{T}&=\bm{F}_{456}(\bm{R}_{t}^{Q,T}\bm{R}_{t}^{s})\end{split}$ (5)
where $\bm{F}_{123}:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$ is the mapping from the $Q$ point position to J1, J2 and J3 joint poses. $\bm{F}_{456}:\mathbb{R}^{3\times 3}\rightarrow\mathbb{R}^{3}$ maps from rotation matrix to J4, J5 and J6 joint poses. $\bm{R}_{t}^{Q}$ is the current orientation of point $Q$.
This decoupling process of the local component’s translation and rotation partially reduces the corresponding number of joints of each local component DoF to less than or equal to three. Although this increases control intuitiveness for human operators, it still suffers from manipulator singularities. In section IV-B, we discuss further amendments that can help avoid manipulator singularities.
Figure 4: Visualization of Singularity Avoidance Motion Mapping. The digital pen in the center indicates the direction where the local component is pointing to. (A) Row angle $d\theta_{t}^{m}$ maps to J1 motion. Notice that under motion mapping (11), the digital pen does not necessaily align with the X-axis at the beginning. (B) Pitch angle $d\phi_{t}^{m}$ maps to J5 motion. (C) Yaw angle $d\psi_{t}^{m}$ maps to J6 motion.
IV-B Singularities in Pieper Kinematics
Following notations in [36], the motion mapping from the configuration space to the operational space has the following form.
$\bm{v}^{e}_{t}=\bm{J}(\bm{q}_{t})\dot{\bm{q}_{t}}$ (6)
An important difference here is that $\bm{v}^{e}_{t}$ now consists of the linear velocity of point $Q$ and the angular velocity of the end-effector $\bm{v}^{e}_{t}=\begin{bmatrix}\dot{\bm{p}}^{Q}_{t}&\bm{\omega}^{e}_{t}\end{bmatrix}$. Since Pieper kinematics structure has six independent joints, the Jacobian can be naturally partitioned into $(3\times 3)$ blocks.
$\bm{J}(\bm{q}_{t})=\begin{bmatrix}\bm{J}_{11}(\bm{q}_{t})&\bm{J}_{12}(\bm{q}_{t})\\ \bm{J}_{21}(\bm{q}_{t})&\bm{J}_{22}(\bm{q}_{t})\end{bmatrix}$ (7)
Manipulator singularities are the joint configurations $\bm{q}$ where the matrix $\bm{J}(\bm{q})$ presents a zero singular value. When $\bm{J}(\bm{q})$ is a square matrix, this is equivalent to having a zero determinant, $\det{\bm{J}(q)}=0$. Since the position $\bm{p}_{t}^{Q}$ is independent of J4-J6, we know that $\bm{J}_{12}(\bm{q}_{t})=\bm{O}_{3}$. As a result, manipulator singularities can be identified from the following equation.
$\det\bm{J}_{11}(\bm{q})\det\bm{J}_{22}(\bm{q})=0$ (8)
Scenarios when $\det\bm{J}_{11}(\bm{q})=0$ is often referred to as arm singularities since it mainly affects the linear velocity and $\det\bm{J}_{22}(\bm{q})=0$ is often referred to as wrist singularities since it mainly affect angular velocities. $\bm{J}_{11}(\bm{q})$ and $\bm{J}_{22}(\bm{q})$ can be written in the form.
$\begin{split}\bm{J}_{11}(\bm{q})&=\begin{bmatrix}((\bm{p}^{Q}-\bm{p}_{1}^{J})\times\bm{\xi}_{1}(\bm{q}))^{T}\\ ((\bm{p}^{Q}-\bm{p}_{2}^{J})\times\bm{\xi}_{2}(\bm{q}))^{T}\\ ((\bm{p}^{Q}-\bm{p}_{3}^{J})\times\bm{\xi}_{3}(\bm{q}))^{T}\end{bmatrix}^{T}\\ \bm{J}_{22}(\bm{q})&=\begin{bmatrix}\bm{\xi}_{4}(\bm{q})&\bm{\xi}_{5}(\bm{q})&\bm{\xi}_{6}(\bm{q})\end{bmatrix}\end{split}$ (9)
where $\bm{\xi}_{i}(\bm{q}),i=1\dots,6$ is the unit vector pointing to the rotational direction of axis $J_{i}$ and $\bm{p}_{i}^{J},i=1\dots 6$ are any random points on axis $J_{i}$. Since J4 is perpendicular to J5 and J5 is perpendicular to J6, $\bm{J}_{22}(\bm{q})=0$ only when J4 is parallel to J6, which happens whenever $q_{5}=0$, as shown in Fig. 3B. One important arm singularity is when $\bm{p}^{Q}-\bm{p}_{2}^{J}$ and $\bm{p}^{Q}-\bm{p}_{2}^{J}$ are parallel, happening when $q_{3}=0,\pm 180^{\circ}$. These are usually called elbow singularities [36].
Arm singularities forms a well-identified set of points in the reachable space $\mathcal{RS}^{Q}$ of point $Q$, so they can be suitably and intuitively avoided by the human operator. Wrist singularities, on the other hand, can be encountered everywhere in the reachable space $\mathcal{RS}^{Q}$, making it overwhelming for the human operator to intentionally avoid them [36]. Thus, in this work, we focus on finding a motion mapping to avoid wrist singularities. In section IV-C, we will discuss this singularity-avoiding motion mapping in detail.
IV-C Singularity-Avoiding Motion Mapping
We first lock J4 at $q_{4}=0$ and generate the Z-Y-X Euler angle leading to $\bm{dR}_{t}^{m}$. Let $\bm{H}$ be the mapping from the a given rotation matrix to the Z-Y-X Euler angles.
$\begin{split}\begin{bmatrix}\theta_{t}^{m}&\phi_{t}^{m}&\psi_{t}^{m}\end{bmatrix}&=\bm{H}(\bm{R}_{t}^{m})\\ \begin{bmatrix}\theta_{0}^{m}&\phi_{0}^{m}&\psi_{0}^{m}\end{bmatrix}&=\bm{H}(\bm{R}_{0}^{m})\end{split}$ (10)
Then, instead of asking the end-effector to follow the rotational translation $\bm{dR}_{t}^{m}$, we directly add the scaled Euler angles to the target J1, J5, and J6 poses.
$\begin{split}q_{1,t}&=q_{1,t-1}+\alpha_{r}(\theta_{t}^{m}-\theta_{0}^{m})\\ q^{\prime}_{5,t}&=q^{\prime}_{5,t-1}+\alpha_{r}(\phi_{t}^{m}-\phi_{0}^{m})\\ q_{6,t}&=q_{6,t-1}+\alpha_{r}(\psi_{t}^{m}-\psi_{0}^{m})\end{split}$ (11)
$q_{1,t}$ and $q_{5,t}$ have been resolved in (11) and $q_{5,t}=q^{\prime}_{5,t}+q_{2,t}+q_{3,t}$ Notice that this is different from computing $\begin{bmatrix}d\theta_{t}^{m}&d\phi_{t}^{m}&d\psi_{t}^{m}\end{bmatrix}=\bm{H}(\bm{dR}_{t}^{m})$ and adding them to $q_{1,t}$, $q^{\prime}_{5,t}$ and $q_{6,t}$.
Without loss of generality, let the local x-direction and y-direction respectively be the normalized master forward vector and master upward vector.
$\begin{split}\bm{fw}_{t}^{m}&=\cos\phi_{t}^{m}\bm{e}_{x}+\sin\phi_{t}^{m}\bm{e}_{y}\\ \bm{up}_{t}^{m}&=\bm{e}_{z}\end{split}$ (12)
where $\bm{e}_{x}$, $\bm{e}_{y}$ and $\bm{e}_{z}$ are respectively X, Y, and Z direction unit vectors in world frame. Similarly, we also define the normalized slave forward vector and slave upward vector.
$\begin{split}\bm{FW}_{t}^{s}&=\cos\phi_{t}^{s}\bm{e}_{x}+\sin\phi_{t}^{s}\bm{e}_{y}\\ \bm{UP}_{t}^{s}&=\bm{e}_{z}\end{split}$ (13)
$q_{2,t}$ and $q_{3,t}$ should satisfy
$\begin{split}(\bm{p}_{t}^{Q}-\bm{p}_{0}^{Q})\cdot\bm{FW}_{t}^{s}&=\alpha_{l}\bm{dp}_{t}^{m}\cdot\bm{fw}_{t}^{m}\\ (\bm{p}_{t}^{Q}-\bm{p}_{0}^{Q})\cdot\bm{UP}_{t}^{s}&=\alpha_{l}\bm{dp}_{t}^{m}\cdot\bm{up}_{t}^{m}\end{split}$ (14)
where $\bm{p}_{t}^{Q}$ depends on $q_{1,t}$, $q_{2,t}$ and $q_{3,t}$ while $q_{1,t}$ is known. (14) can be easily solved using trigonometry, and this process is singularity-free for most robot configurations. Finally, $q_{5,t}$ is obtained by
$q_{5,t}=q^{\prime}_{5,t}-q_{2}-q_{3}$ (15)
Since (11) depends on the local component’s relative motion, although mapping $\bm{H}$ has singularity when $\theta_{0}^{m}$ or $\theta_{t}^{m}=\pm 90^{\circ}$, the user can easily avoid this point via resetting the starting master pose $\bm{R}_{0}^{s}$. The simple additions in (11) and (15) indicates that as long as $\theta_{t}^{m}$ or $\theta_{0}^{m}$ in the local component are kept away from ±$\pm 90^{\circ}$, the slave system motion will always be smooth even in wrist singularities.
An important property of the motion mapping in (10) and (11) is that if the starting pose of the local component $\bm{R}_{0}^{m}$ aligns with the slave end-effector $\bm{R}_{0}^{s}$, then they will be pointing at the same direction at all time, $\bm{fw}_{t}^{m}=\bm{FW}_{t}^{s}$, if $\alpha_{r}=1$.
However, this singularity avoidance mapping comes with a price, that being the lose of the J4 DoF. $q_{4}$ will be maintained at zero throughout local control. Unfortunately, as argued in section IV-B, if J4 is released and $q_{5}$ is very close to zero, a random orientation-preserved linear motion with small amplitude of the local component will cause a large movement of the slave system. Therefore, the authors believe that developing a local motion mapping that preserves all 6 DoFs of the slave end-effector while maintaining smooth transitions at singularities requires new designs on the slave side. This will be discussed in detail in section VII.
V Operational Jacobian in Teleoperation
In this part, we give a more general and mathematically rigorous analysis of why the singularity-avoiding mapping can avoid wrist singularities.
Denote the generalized position of the master device with vectors $\bm{P}$, which can either be joint poses in the global component or the spatial position defined in the $SE(3)$ set. $\bm{q}$ is the generalized position of the slave system, which usually consists of the slave joint poses. We spot that relation between the master device motion $\dot{\bm{P}}\in\mathbb{R}^{n}$ and slave side motion $\dot{\bm{q}}\in\mathbb{R}^{m}$ often has the following form
$\mathcal{F}(\bm{P})\dot{\bm{P}}=\bm{J}_{o}(\bm{P},\bm{q})\dot{\bm{q}}$ (16)
where $\mathcal{F}(\bm{P})\in\mathbb{R}^{\min(m,n)\times n}$ decreases dimension to handle the case when $n>m$. We name $\bm{J}_{o}(\bm{P},\bm{q}):\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{m\times m}$ the operational Jacobian, which can be used to indicate the smoothness of the motion mapping. The configuration that causes $\bm{J}_{o}$ to have a zero singular value is called operational singularity, meaning that certain local component motion will cause indefinite joint motion in the slave system. For our local component, which is a haptic device, there is $\dot{\bm{P}}=\begin{bmatrix}\bm{v}^{o}&\bm{\omega}^{o}\end{bmatrix}$, where $\bm{v}^{o}$ and $\bm{\omega}^{o}$ are the stylus linear and angular velocities.
In direct motion mapping as shown in Section III, there are relations $v_{i}^{o}=\alpha_{l}v_{i}$ and $\omega_{i}^{o}=\alpha_{r}\omega_{i}$ from (2), indicating $\mathcal{F}^{DM}(\bm{P})=\bm{I}_{6}$ and the following relation between the operational Jacobian and manipulator Jacobian.
$\bm{J}^{DM}_{o}(\bm{P},\bm{q}_{t})=\begin{bmatrix}\alpha_{l}\bm{I}_{3}&\bm{O}\\ \bm{O}&\alpha_{r}\bm{I}_{3}\end{bmatrix}\bm{J}_{e}(\bm{q}_{t})$ (17)
Therefore, $\bm{J}^{DM}_{o}$ inherits all singularities of the manipulator Jacobian. Since $\bm{J}^{DM}_{o}$ depends only on $\bm{q}_{t}$, it is infeasible to achieve smooth transition around singularities via adjusting the starting pose $\bm{p}_{0}^{m}$.
On the other hand, for the singularity avoidance mapping in section IV-C, since $q_{4}$ is fixed at 00, we have $m=5$ and $n=6$. Its operational Jacobian and projector can be written as the following
$\begin{split}\bm{J}^{SA}_{o}(\bm{P},\bm{q})&=\begin{bmatrix}0&A_{11}&A_{12}&0&0\\ 0&A_{21}&A_{22}&0&0\\ 1&0&0&0&\cos{\theta_{t}^{m}}\\ 0&0&0&\sin{\phi_{t}^{m}}&\sin{\theta_{t}^{m}}\cos{\phi_{t}^{m}}\\ 0&0&0&-\cos{\phi_{t}^{m}}&\sin{\theta_{t}^{m}}\sin{\phi_{t}^{m}}\end{bmatrix}W\\ \mathcal{F}^{SA}(\bm{P})&=\begin{bmatrix}\cos{\phi_{t}^{m}}&\sin{\phi_{t}^{m}}&0&0&0&0\\ 0&0&1&0&0&0\\ 0&0&0&1&0&0\\ 0&0&0&0&1&0\\ 0&0&0&0&0&1\end{bmatrix}\end{split}$
$W$ is a constant matrix generated from $q_{5}^{\prime}=q_{5}+q_{2}+q_{3}$.
$W=\begin{bmatrix}1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\\ 0&1&1&1&0\\ 0&0&0&0&1\end{bmatrix}$ (18)
Letting the distance from J2 to J3 and the distance from J3 to $Q$ to be $l_{2}$ and $l_{3}$ respectively, we have
$\begin{split}A_{11}&=l_{2}\sin{q_{2}}-l_{3}\sin(q_{2}+q_{3})\\ A_{12}&=-l_{3}\sin(q_{2}+q_{3})\\ A_{21}&=l_{2}\cos{q_{2}}-l_{3}\cos(q_{2}+q_{3})\\ A_{22}&=-l_{3}\cos(q_{2}+q_{3})\end{split}$ (19)
Since $W$ and $\mathcal{F}^{SA}(\bm{P})$ always have nonzero singular values, identifying the singularities in (V) is equivalent to identifying singularties of the following two submatrices.
$\bm{A}=\begin{bmatrix}A_{11}&A_{12}\\ A_{21}&A_{22}\\ \end{bmatrix}\quad\bm{B}=\begin{bmatrix}1&0&\cos{\theta_{t}^{m}}\\ 0&\sin{\phi_{t}^{m}}&\cos{\theta_{t}^{m}}\cos{\phi_{t}^{m}}\\ 0&-\cos{\phi_{t}^{m}}&\cos{\theta_{t}^{m}}\sin{\phi_{t}^{m}}\end{bmatrix}$
$\bm{A}$ has zero singular value at the elbow singularity, which can be easily and intuitively avoided by the human operators. $\bm{B}$ is singular when $\theta_{t}^{m}=\pm 90^{\circ}$. Interestingly, $\theta_{t}^{m}$ here is the pitch angle of the local component instead of the slave end-effector. This means one can always reset the local component starting orientation $\bm{R}_{0}^{m}$ to avoid $\theta_{t}^{m}$ from reaching ±$\pm 90^{\circ}$. Therefore, singularities in $\bm{B}$ is independent of slave joint poses $\bm{q}$, showing that wrist singularities is no longer impeding the teleoperation performance.
Motion mapping (16) covers a much wider range of teleoperation systems and does not require the master and slave systems to have the same DoF. Therefore, the idea of using operational Jacobians in analyzing the teleoperation system’s robustness against manipulator singularities can also be generalized to teleoperation systems with different kinematic structures.
Figure 5: Hardware setup of the Global–Local (G–L) teleoperation interface. (A), (B): Global teleoperation subsystem using a 3-DOF joint-link chain as the master device. (C), (D): Local teleoperation subsystem using a commercial haptic device as the master component.

### v3

@@ II-A Temporal Decoupling
In the temporal decoupling scheme, the global component is realized as a scaled replica of the slave system, a common approach in teleoperation [18, 35]. The local component utilizes a haptic device for precise tracking of the operator’s fingertip position and orientation. Activating the global component places the slave system in joint control mode, replicating the scaled replica’s joint positions within the limits of the motor capacities. Conversely, activating the local component switches the slave system to end-effector Cartesian control, where it tracks the motion of the haptic device. Fig. 2 shows a schematic diagram of temporal decoupling.
To facilitate precise control of the end-effector, the slave system is designed to track only a portion of the local component motion. Every time the local component is activated, its current position $p_{0}$ and orientation $q_{0}$ will be re-initialized as the origin. Given a linear scaling factor $\alpha_{l}\in(0,1]$, the slave system will move its end-effector position towards the scaled linear displacement of the haptic device, $\alpha_{l}(p-p_{0})$. For rotational motions, assume the orientation of the local component has moved from $q_{0}$ to $q$, we convert the rotational displacement $q_{0}^{-1}q$ into axis-angle representation $(\vec{v},\theta)$, where $\vec{v}$ is the rotational axis and $\theta$ is the rotational angle. Assume the slave system’s end-effector currently has orientation $q_{e}$, given a rotational scaling factor $\alpha_{r}\in(0,1]$, the slave system will target to change its end-effector orientation via rotating around axis $\vec{v}$ by angle $\alpha_{r}\theta$, starting from $q_{e}$.
Figure 2: A schematic diagram of temporal decoupling. The global component is a scaled replica of the slave robot and the local component is a haptic device. The two step pedals allows the operator to rapidly switch between the local and global component. If the local component is activated, the global component will temporary turn into a slave system that follows the actual slave system to avoid joint position mismatch when switching back to the global component.
This design presents a challenge during the transition from the local to global component. A mismatch can arise between the pose of the scaled replica (global component) and the actual slave system configuration, potentially causing an abrupt motion in the slave upon returning to global control. To mitigate this, the global component assumes a ”slave” role during local component activation, mirroring the joint motions of the true slave system. This ensures pose synchronization between the global component and the slave, facilitating a smooth transition when switching back to global controls.
It is worth noting that there is an overlap between the DOFs that the global and local components are controlling, so they cannot be activated at the same time. Using devices like foot pedals allows the operator to conveniently switch between the global and local components.
II-B Spatial Decoupling
Figure 3: A schematic diagram of spatial decoupling. The global component is a scaled replcia of the first N3N-3 DOFs, while the local component consists of two IMU sensors to measure human operator’s wrist rotation and map it back to the last three joints’s target positions. Since it eliminates all mechanical constraints around the operator’s hand, it allows mounting complicated end-effector controllers, such as an exoskeleton structure for controlling a dexterous hand.
Spatial decoupling leverages a common characteristic of serially-linked industrial robots in that the final three joints primarily influence end-effector orientation with less impact on the overall robot configuration, while the remaining joints predominantly determine end-effector position and potential collisions. This principle is applied to an $N$-DOF industrial robot (e.g., KUKA or FANUC) by dividing control into two components. The global component replicates the first N3N-3 joints, providing gross positioning, while a local component, consisting of a dual-IMU system, mounted on the forearm and the hand of the human operator. measures the relative rotation $R$ between the human wrist and hand to control the final three joints. This approach often simplifies inverse kinematics to finding X-Y-Z or X-Y-X Euler angles based on $R$. Fig. 3 shows a schematic diagram of spatial decoupling.
This decoupling architecture allows simultaneous activation of both components. The operator can manipulate the global component with one hand and the local component with the other. While seemingly adding complexity, this approach offers two key advantages. First, the dual-IMU design in the local component eliminates mechanical constraints around the operator’s hand, allowing integration of additional devices to control more sophisticated end-effectors. Second, replicating the first N3N-3 joints enables full control of the slave robot’s redundant DOFs, maximizing workspace utilization. Existing teleoperation systems often prioritize one of these advantages. Systems focusing on precise hand pose tracking [3, 2, 31] may have limited control over the full robot workspace. Conversely, systems emphasizing workspace utilization [18, 35, 1] may impose constraints on the operator’s hand, potentially hindering dexterity and the use of additional tools. Spatial decoupling combines both advantages, as demonstrated in the experiment section where a glove-shaped device is mounted on the local component to teleoperate a multi-fingered hand.
Figure 4: Physical realization of temporal decoupling. (A), (B) Teleoperation with global component. (C), (D) Teleoperation with local component.
Figure 5: Physical realization of spatial decoupling. (A) the operator moves the global and local components simultaneously. (B) The operator moves the local component only to manipulate the last three joints of the Flexiv robot arm.
III Implementation Details
Figure 6: The two-finger exoskeleton design in the spatial decomposition framework. The thumb of the Inspire hand follows the human operator’s thumb, and the index, middle, ring, and pinky fingers follow the human operator’s index. (B) shows how the two MTI-3 sensors are mounted on the operator’s arm for wrist rotation measurement.
Figure 4 and 5 depicts the two teleoperation systems developed to implement the proposed temporal and spatial decoupling framework. The dual-arm teleoperation system, designed according to temporal decoupling, utilizes two Piper Agile X Limit robotic arms as the slave system. For gross manipulation, the global component employs Dynamixel servomotors, while the local component integrates two Touch X haptic devices for enhanced fine control. Each Piper arm, a 6-DOF lightweight manipulator, is equipped with a two-fingered gripper for position-controlled grasping. The master system’s global component comprises six Dynamixel servomotors: two XL430 motors for the base joints and four XL330-M288 motors for the remaining degrees of freedom. The Touch X devices, serving as the local component, are 6-DOF haptic interfaces that track the human operator’s finger tip position and orientation precisely. Similar to the single-arm system, all devices are managed and coordinated through a multi-process LabVIEW program.